# Scheduling problems of aircraft on a same runway and dual runways

### v7

@@ IV Takeoff scheduling problem on a single runway
Consider a takeoff sequence $\Phi_{a}=\langle\phi_{1},\phi_{2}\rangle$, where $\phi_{1}=\langle Tcf_{1},Tcf_{2},\cdots,Tcf_{n_{1}}\rangle$, $\phi_{2}=\langle Tcf_{n_{1}+1},$ $Tcf_{{n_{1}}+2},\cdots,Tcf_{n_{2}}\rangle$, $\phi_{1}$ and $\phi_{2}$ are both class-monotonically-decreasing sequences, $cl_{n_{1}}<cl_{n_{1}+1}$, and $n_{1},n_{2}$ are two positive integers. Suppose that Assumptions 6-10 hold for each takeoff aircraft sequence. Merge the aircraft sequences $\phi_{1}$ and $\phi_{2}$ to form a new class-monotonically-decreasing sequence $\Phi_{b}$.
@@
Consider a takeoff aircraft sequence $\Phi_{a}=\langle\phi_{1},\phi_{2}\rangle$, where $\phi_{1}=\langle Tcf_{1},Tcf_{2},\cdots,Tcf_{s_{1}}\rangle$, $\phi_{2}=\langle Tcf_{s_{1}+1},Tcf_{s_{1}+2},\cdots,Tcf_{n}\rangle$, both of the sequences $\phi_{1}$ and $\phi_{2}$ are class-monotonically-decreasing sequences, $cl_{s_{1}}<cl_{s_{1}+1}$, and $s_{1}<n-1$ is a positive integer. Suppose that $\phi_{2}$ contains $s_{2}$ aircraft of class $h_{1}$, $Tcf_{i}\in\phi_{2}$, and $cl_{i}=h_{1}<cl_{s_{1}}$. Suppose that Assumptions 6-10 hold for each aircraft sequence. Generate a new sequence $\Phi_{b}$ by moving the aircraft $Tcf_{i}$ to be between $Tcf_{s_{1}}$ and $Tcf_{s_{1}+1}$. The following statements hold.
@@
Consider a takeoff sequence $\Phi_{a}=\langle\phi_{1},\phi_{2},\cdots,\phi_{s}\rangle$ for a positive integer $s$, where $\phi_{i}=\langle Tcf_{i1},Tcf_{i2},\cdots,$ $Tcf_{ic_{i}}\rangle$ is a class-monotonically-decreasing sequence for some positive integer $c_{i}$ and all $i\in\{1,2,\cdots,s\}$, and $cl_{jc_{j}}<cl_{(j+1)1}$ for all $j\in\{1,2,\cdots,s-1\}$. Suppose that Assumptions 6-10 hold for each takeoff aircraft sequence. Merge the aircraft sequences $\Phi_{a}$ to form a new class-monotonically-decreasing sequence $\Phi_{b}$. Then, $F(\Phi_{a})=F(\Phi_{b})+D_{d1}-D_{cl_{sc_{s}}1}-\mathrm{sgn}(\sum_{i=1}^{s}\mathrm{Exs}_{\rho_{1}}(\phi_{i}))[\sum_{i=1}^{s}\mathrm{Exs}_{\rho_{1}}(\phi_{i})-1]T_{0}/3-\mathrm{sgn}(\sum_{i=1}^{s}\mathrm{Exs}_{\rho_{2}}(\phi_{i}))[\sum_{i=1}^{s}\mathrm{Exs}_{\rho_{2}}(\phi_{i})-1]T_{0}/3+\sum_{i=1}^{s-1}[D_{cl_{ic_{i}}cl_{(i+1)1}}-D_{cl_{ic_{i}}1}]$, where $d$ denotes the class of the last aircraft of $\Phi_{b}$.
@@ V Scheduling problem of landing and takeoff aircraft on a same runway
(Path) Consider a sequence $\phi=\langle Tcf_{1},Tcf_{2},\cdots,Tcf_{n}\rangle$. For any given aircraft $Tcf_{i}$ and $Tcf_{j}$, if there exists an aircraft subsequence $\langle Tcf_{i}^{0},Tcf_{i}^{1},\cdots,Tcf_{i}^{\rho}\rangle$ for some positive integer $\rho>0$ such that $Tcf_{i}^{0}=Tcf_{i}$, $Tcf_{i}^{\rho}=Tcf_{j}$ and each aircraft $Tcf^{h}_{i}$ is relevant to aircraft $Tcf^{h-1}_{i}$, $i=1,2,\cdots,\rho$, then the sequence $\langle Tcf_{i}^{0},Tcf_{i}^{1},\cdots,Tcf_{i}^{\rho}\rangle$ is said to be a path from the aircraft $Tcf_{j}$ to the aircraft $Tcf_{i}$. It is assumed by default that each aircraft has a path to itself.
@@
Consider an aircraft sequence $\phi_{1}=\langle\phi_{11},\phi_{12}\rangle$, where $\phi_{11}=\langle Tcf_{1},Tcf_{2},Tcf_{3},Tcf_{4},Tcf_{5},Tcf_{6}\rangle$, all the aircraft of $\phi_{11}$ are takeoff (landing) aircraft, $\phi_{12}=\langle Tcf_{7},Tcf_{8},Tcf_{9},Tcf_{10}\rangle$, all the aircraft of $\phi_{12}$ are landing (takeoff) aircraft. Suppose that Assumption 14 holds.
@@
Consider an aircraft sequence $\phi^{1}_{0}=\phi^{2}_{0}=\langle Tcf_{i},Tcf_{i+1},Tcf_{i+2}\rangle$, where the aircraft $Tcf_{i}$ and $Tcf_{i+2}$ are both takeoff (landing) aircraft, the aircraft $Tcf_{i+1}$ is a landing (takeoff) aircraft, and the separation times between the aircraft in $\phi^{1}_{0}$ and $\phi^{2}_{0}$ are different. Suppose that in $\phi_{0}^{1}$ and $\phi_{0}^{2}$, the aircraft $Tcf_{i+2}$ is relevant to either the aircraft $Tcf_{i}$ or the aircraft $Tcf_{i+1}$. When $S_{i(i+1)}(\phi_{0}^{1})>S_{i(i+1)}(\phi_{0}^{2})$, then $S_{(i+1)(i+2)}(\phi_{0}^{1})=S_{(i+1)(i+2)}(\phi_{0}^{2})$ or $S_{(i+1)(i+2)}(\phi_{0}^{1})<S_{(i+1)(i+2)}(\phi_{0}^{2})$. Further, when $S_{i(i+1)}(\phi_{0}^{1})>S_{i(i+1)}(\phi_{0}^{2})$ and $S_{(i+1)(i+2)}(\phi_{0}^{1})<S_{(i+1)(i+2)}(\phi_{0}^{2})$, $S_{i(i+2)}(\phi_{0}^{2})=Y_{i(i+2)}$, i.e., in $\phi_{0}^{2}$, the aircraft $Tcf_{i+2}$ is relevant to the aircraft $Tcf_{i}$.
@@ VI Scheduling problem of landing and takeoff aircraft on dual runways
Suppose that an aircraft $Tcf_{i}$ and a trailing aircraft $Tcf_{j}$ consecutively take off from or land on dual runways whose spacing is no larger than $760$ $m$.
(1) Suppose that $D_{P}=T_{0}$ and $P_{D}=0$, when the two aircraft have different operation tasks.
(2) Suppose that the minimum separation time is equal to that on the same runway, when the two aircraft $Tcf_{i}$ and $Tcf_{j}$ are both landing/takeoff aircraft.
@@
Suppose that the orders of all aircraft in $\Phi$ is fixed. Suppose that $Tf_{k}=[t_{0},+\infty]$ for all $k$. Under Assumptions 1-4, 6-9 and 15, the following statements hold.
@@
Suppose that for each aircraft $Tcf_{i}\in\{2,3,\cdots,n+m\}$ in $\Phi$, the aircraft $Tcf_{i}$ is relevant to the aircraft $Tcf_{k}$ for some $k\in\{Tcf_{i-4},Tcf_{i-3},Tcf_{i-2},Tcf_{i-1}\}$. If aircraft $Tcf_{i}$ and $Tcf_{i-1}$ have different (takeoff/landing) operation tasks and aircraft $Tcf_{i}$ is relevant to $Tcf_{i-1}$ for $i\in\{2,3,\cdots,n+m\}$, there is at least an aircraft $Tcf_{j}\in\{Tcf_{1},Tcf_{2},\cdots,Tcf_{k-1}\}$ such that $S_{ji}(\Phi)-T_{0}<Y_{ji}$.
@@
Suppose that $Tf_{j}=[t_{0},+\infty]$ for all $j$ and Assumptions 1-4, 6-9 and 15, 16 hold.
@@
(1) Suppose that $Tcf_{i_{1}-1}\in\phi_{0}$, $Tcf_{i_{1}}\in\phi_{1}$, for the sequence $\phi_{0}$, there is a path from its last aircraft $Tcf_{i_{1}-1}$ to its first aircraft, and for the sequence $\phi_{1}$, there is a path from its second last aircraft to its first aircraft. If $Tcf_{i_{1}}$ is relevant to only $Tcf_{i_{1}-1}$, it is said that the sequence $\phi$ is a T-block subsequence of $\Phi$, and $(|\phi_{1}|-1)/(|\phi_{0}|-1)$ and $S_{i_{1}}(\phi)-S_{i_{2}}(\phi)-Y_{i_{2}i_{1}}$ are the length and the takeoff time increment of T-block subsequence, where $Tcf_{i_{2}}$ is the second last takeoff aircraft in $\phi$. Let $\langle Tcf_{k_{0}},Tcf_{k_{1}},\cdots,Tcf_{k_{s}}\rangle$ denote the takeoff sequence in the T-block subsequence $\phi$. Consider a subsequence $\bar{\phi}$ such that the landing times of the landing aircraft in $\bar{\phi}$ are consistent with those in $\phi$ and the takeoff times of the takeoff aircraft satisfy that $S_{k_{0}}(\bar{\phi})-S_{k_{0}}(\phi)=S_{k_{1}}(\bar{\phi})-S_{k_{1}}(\phi)=\cdots=S_{k_{s-1}}(\bar{\phi})-S_{k_{s-1}}(\phi)=\mu$ and $S_{k_{s}}(\bar{\phi})=S_{k_{s}}(\phi)$ for some constant $\mu\geq 0$. Suppose that for a given constant $\mu_{0}$, if $0\leq\mu\leq\mu_{0}$, the minimum separation requirements can be satisfied in $\bar{\phi}$ and if $\mu>\mu_{0}$, the minimum separation requirements cannot be satisfied in $\bar{\phi}$. Then the constant $\mu_{0}$ is said to be the initial redundant time of the T-block subsequence $\phi$.
@@
(2) Suppose that $Tcf_{i_{1}-1}\in\phi_{1}$, $Tcf_{i_{1}}\in\phi_{0}$, for the sequence $\phi_{1}$, there is a path from its last aircraft $Tcf_{i_{1}-1}$ to its first aircraft, and for the sequence $\phi_{0}$, there is a path from the second last aircraft to the first aircraft in $\phi_{0}$. If $Tcf_{i_{1}}$ is relevant to only $Tcf_{i_{1}-1}$, it is said that the sequence $\phi$ is a D-block subsequence of $\Phi$, and $(|\phi_{0}|-1)/(|\phi_{1}|-1)$ and $S_{i_{1}}(\phi)-S_{i_{2}}(\phi)-Y_{i_{2}i_{1}}$ are the length and the landing time increment of D-block subsequence, where $Tcf_{i_{2}}$ is the second last landing aircraft in $\phi$. Let $\langle Tcf_{k_{0}},Tcf_{k_{1}},\cdots,Tcf_{k_{s}}\rangle$ denote the landing sequence in the D-block subsequence $\phi$. Consider a subsequence $\bar{\phi}$ such that the takeoff times of the takeoff aircraft in $\bar{\phi}$ are consistent with those in $\phi$ and $S_{k_{0}}(\bar{\phi})-S_{k_{0}}(\phi)=S_{k_{1}}(\bar{\phi})-S_{k_{1}}(\phi)=\cdots=S_{k_{s-1}}(\bar{\phi})-S_{k_{s-1}}(\phi)=\mu$ and $S_{k_{s}}(\bar{\phi})=S_{k_{s}}(\phi)$ for some constant $\mu\geq 0$. Suppose that for a given constant $\mu_{0}$, if $0\leq\mu\leq\mu_{0}$, the minimum separation requirements can be satisfied in $\bar{\phi}$ and if $\mu>\mu_{0}$, the minimum separation requirements cannot be satisfied in $\bar{\phi}$. Then the constant $\mu_{0}$ is said to be the initial redundant time of the D-block subsequence $\phi$.
@@
From Tables I-IV, it can be seen that once the separation times between the landing/takeoff aircraft are fixed, in most of the scenarios, each of the takeoff/landing aircraft is relevant to an aircraft with different operation task, which form block subsequences of length $1/(*)$. There are only a very small number of scenarios marked in red such that each of the takeoff/landing aircraft is relevant to only an aircraft with the same operation task and block subsequences are not formed. For example, consider an optimal matching problem between two sequences, $\phi_{0}=\langle Tcf_{1},Tcf_{2},Tcf_{3}\rangle$ and $\phi_{1}=\langle Tcf_{4},Tcf_{5},Tcf_{6}\rangle$, where $S_{1}(\phi_{0})=S_{4}(\phi_{1})$, $Y_{12}=Y_{23}=90$, $Y_{45}=60$, and $Y_{56}=80$. If we let $S_{12}(\phi_{0})=S_{23}(\phi_{0})=90$, $S_{45}(\phi_{1})$ and $S_{56}(\phi_{1})$ should be taken as $S_{45}(\phi_{1})=S_{56}(\phi_{1})=90$, whereas if we let $S_{45}(\phi_{1})=60$ and $S_{56}(\phi_{1})=80$, $S_{12}(\phi_{0})$ should be taken as $S_{12}(\phi_{0})=140$, and $S_{23}(\phi_{0})$ can be taken as $S_{23}(\phi_{0})=90$ if there are no other aircraft. In this example, it can be seen that the matching between a landing sequence and a takeoff sequence is not very complicated and the number of all possible block subsequences is $2$.
From Tables V-VIII, it can be seen that the range of the time increments of block subsequences is not large as well, in particular when the takeoff time increments are considered. As a matter of fact, based on the calculations for more general scenarios including those listed in Tables I-VIII, it is found that if the number of breakpoint aircraft in a block subsequence is fixed, the number of all possible block subsequences of different class combinations and the range of the takeoff/landing time increments of block subsequences are both not large according to the minimum separation time standards at Heathrow Airport and for the RECAT-EU system in the absence of resident-point aircraft. This property can be fully explored to find the optimal sequence when a landing sequence and a takeoff sequence are matched.

### v8

@@ III Landing scheduling problem on a single runway
When the aircraft are free of time window constraints, it is assumed by default that the operation time of the first aircraft of a sequence is $t_{0}$ if no otherwise specified.
@@ IV Takeoff scheduling problem on a single runway
Consider a takeoff sequence $\Phi_{a}=\langle\phi_{1},\phi_{2}\rangle$, where $\phi_{1}=\langle Tcf_{1},Tcf_{2},\cdots,Tcf_{n_{1}}\rangle$, $\phi_{2}=\langle Tcf_{n_{1}+1},$ $Tcf_{{n_{1}}+2},\cdots,Tcf_{n_{2}}\rangle$, $\phi_{1}$ and $\phi_{2}$ are both class-monotonically-decreasing sequences, $cl_{n_{1}}<cl_{n_{1}+1}$, and $n_{1},n_{2}$ are two positive integers. Merge the aircraft sequences $\phi_{1}$ and $\phi_{2}$ to form a new class-monotonically-decreasing sequence $\Phi_{b}$. Suppose that Assumptions 6-10 hold for each takeoff aircraft sequence. The following statements hold.
@@
Consider a takeoff aircraft sequence $\Phi_{a}=\langle\phi_{1},\phi_{2}\rangle$, where $,\phi_{1}=\langle Tcf_{1},Tcf_{2},$$\cdots,Tcf_{s_{1}}\rangle$, $\phi_{2}=\langle Tcf_{s_{1}+1},Tcf_{s_{1}+2},\cdots,Tcf_{n}\rangle$, both of the sequences $\phi_{1}$ and $\phi_{2}$ are class-monotonically-decreasing sequences, $cl_{s_{1}}<cl_{s_{1}+1}$, and $s_{1}<n-1$ is a positive integer. Suppose that $\phi_{2}$ contains $s_{2}$ aircraft of class $h_{1}$, $Tcf_{i}\in\phi_{2}$, and $cl_{i}=h_{1}<cl_{s_{1}}$. Generate a new sequence $\Phi_{b}$ by moving the aircraft $Tcf_{i}$ to be between $Tcf_{s_{1}}$ and $Tcf_{s_{1}+1}$. Suppose that Assumptions 6-10 hold for each aircraft sequence. The following statements hold.
@@
Consider a takeoff sequence $\Phi_{a}=\langle\phi_{1},\phi_{2},\cdots,\phi_{s}\rangle$ for a positive integer $s$, where $\phi_{i}=\langle Tcf_{i1},Tcf_{i2},\cdots,$ $Tcf_{ic_{i}}\rangle$ is a class-monotonically-decreasing sequence for some positive integer $c_{i}$ and all $i\in\{1,2,\cdots,s\}$, and $cl_{jc_{j}}<cl_{(j+1)1}$ for all $j\in\{1,2,\cdots,s-1\}$. Merge the aircraft sequences $\Phi_{a}$ to form a new class-monotonically-decreasing sequence $\Phi_{b}$. Suppose that Assumptions 6-10 hold for each sequence. Then, $F(\Phi_{a})=F(\Phi_{b})+D_{d1}-D_{cl_{sc_{s}}1}-\mathrm{sgn}(\sum_{i=1}^{s}\mathrm{Exs}_{\rho_{1}}(\phi_{i}))[\sum_{i=1}^{s}\mathrm{Exs}_{\rho_{1}}(\phi_{i})-1]T_{0}/3-\mathrm{sgn}(\sum_{i=1}^{s}\mathrm{Exs}_{\rho_{2}}(\phi_{i}))[\sum_{i=1}^{s}\mathrm{Exs}_{\rho_{2}}(\phi_{i})-1]T_{0}/3+\sum_{i=1}^{s-1}[D_{cl_{ic_{i}}cl_{(i+1)1}}-D_{cl_{ic_{i}}1}]$, where $d$ denotes the class of the last aircraft of $\Phi_{b}$.
@@ V Scheduling problem of landing and takeoff aircraft on a same runway
(Path) Consider a sequence $\phi=\langle Tcf_{1},Tcf_{2},\cdots,Tcf_{n}\rangle$. For any given aircraft $Tcf_{i}$ and $Tcf_{j}$, if there exists an aircraft subsequence $\langle Tcf_{i}^{0},Tcf_{i}^{1},\cdots,Tcf_{i}^{\rho}\rangle$ for some positive integer $\rho>0$ such that $Tcf_{i}^{0}=Tcf_{i}$, $Tcf_{i}^{\rho}=Tcf_{j}$ and each aircraft $Tcf^{h}_{i}$ is relevant to aircraft $Tcf^{h-1}_{i}$, $h=1,2,\cdots,\rho$, then the sequence $\langle Tcf_{i}^{0},Tcf_{i}^{1},\cdots,Tcf_{i}^{\rho}\rangle$ is said to be a path from the aircraft $Tcf_{j}$ to the aircraft $Tcf_{i}$. It is assumed by default that each aircraft has a path to itself.
@@
Consider an aircraft sequence $\phi_{1}=\langle\phi_{11},\phi_{12}\rangle$, where $\phi_{11}=\langle Tcf_{1},Tcf_{2},Tcf_{3},Tcf_{4},Tcf_{5},Tcf_{6}\rangle$, all the aircraft of $\phi_{11}$ are takeoff (landing) aircraft, $\phi_{12}=\langle Tcf_{7},Tcf_{8},Tcf_{9},Tcf_{10}\rangle$, all the aircraft of $\phi_{12}$ are landing (takeoff) aircraft. Suppose that Assumption 14 holds for each sequence.
@@
Consider an aircraft sequence $\phi^{1}_{0}=\phi^{2}_{0}=\langle Tcf_{i},Tcf_{i+1},Tcf_{i+2}\rangle$, where the aircraft $Tcf_{i}$ and $Tcf_{i+2}$ are both takeoff (landing) aircraft, the aircraft $Tcf_{i+1}$ is a landing (takeoff) aircraft, and the separation times between the aircraft in $\phi^{1}_{0}$ and $\phi^{2}_{0}$ are different. Suppose that in $\phi_{0}^{1}$ and $\phi_{0}^{2}$, $S_{i}(\phi^{1}_{0})=S_{i}(\phi^{2}_{0})$ and the aircraft $Tcf_{i+2}$ is relevant to either the aircraft $Tcf_{i}$ or the aircraft $Tcf_{i+1}$. When $S_{i(i+1)}(\phi_{0}^{1})>S_{i(i+1)}(\phi_{0}^{2})$, then $S_{(i+1)(i+2)}(\phi_{0}^{1})=S_{(i+1)(i+2)}(\phi_{0}^{2})$ or $S_{(i+1)(i+2)}(\phi_{0}^{1})<S_{(i+1)(i+2)}(\phi_{0}^{2})$. Further, when $S_{i(i+1)}(\phi_{0}^{1})>S_{i(i+1)}(\phi_{0}^{2})$ and $S_{(i+1)(i+2)}(\phi_{0}^{1})<S_{(i+1)(i+2)}(\phi_{0}^{2})$, $S_{i(i+2)}(\phi_{0}^{2})=Y_{i(i+2)}$, i.e., in $\phi_{0}^{2}$, the aircraft $Tcf_{i+2}$ is relevant to the aircraft $Tcf_{i}$.
@@ VI Scheduling problem of landing and takeoff aircraft on dual runways
Suppose that $D_{P}=T_{0}$, $P_{D}=0$, and for any two aircraft $Tcf_{i}$ and $Tcf_{j}$ with the same operation tasks, $Y_{ij}=T_{cl_{i}cl_{j}}$ or $Y_{ij}=D_{cl_{i}cl_{j}}$.
@@
Consider the sequence $\Phi$. Suppose that the orders of all aircraft in $\Phi$ is fixed and $Tf_{k}=[t_{0},+\infty]$ for all $k$. Under Assumptions 1-4, 6-9 and 15, the following statements hold.
@@
Suppose that for each aircraft $Tcf_{i}\in\{2,3,\cdots,n+m\}$ in $\Phi$, the aircraft $Tcf_{i}$ is relevant to an aircraft $Tcf_{k}$ for $Tcf_{k}\in\{Tcf_{i-4},Tcf_{i-3},Tcf_{i-2},Tcf_{i-1}\}$. If aircraft $Tcf_{i}$ and $Tcf_{i-1}$ have different (takeoff/landing) operation tasks and aircraft $Tcf_{i}$ is relevant to $Tcf_{i-1}$ for $i\in\{2,3,\cdots,n+m\}$, there is at least an aircraft $Tcf_{j}\in\{Tcf_{1},Tcf_{2},\cdots,Tcf_{k-1}\}$ such that $S_{ji}(\Phi)-T_{0}<Y_{ji}$.
@@
Consider the sequence $\Phi$. Suppose that $Tf_{j}=[t_{0},+\infty]$ for all $j$ and Assumptions 1-4, 6-9 and 15, 16 hold.
@@
(1) Suppose that $Tcf_{i_{1}-1}\in\phi_{0}$, $Tcf_{i_{1}}\in\phi_{1}$, for the sequence $\phi_{0}$, there is a path from its last aircraft $Tcf_{i_{1}-1}$ to its first aircraft, and for the sequence $\phi_{1}$, there is a path from its second last aircraft to its first aircraft. If $Tcf_{i_{1}}$ is relevant to only $Tcf_{i_{1}-1}$, it is said that the sequence $\phi$ is a T-block subsequence of $\Phi$, and $(|\phi_{1}|-1)/(|\phi_{0}|-1)$, $S_{i_{1}}(\phi)-S_{i_{0}}(\phi)$ and $S_{i_{1}}(\phi)-S_{i_{2}}(\phi)-Y_{i_{2}i_{1}}$ are the capacity, the time length and the takeoff time increment of T-block subsequence, where $Tcf_{i_{2}}$ is the second last takeoff aircraft in $\phi$. Let $\langle Tcf_{k_{0}},Tcf_{k_{1}},\cdots,Tcf_{k_{s}}\rangle$ denote the takeoff sequence in the T-block subsequence $\phi$. Consider a subsequence $\bar{\phi}$ such that the landing times of the landing aircraft in $\bar{\phi}$ are consistent with those in $\phi$ and the takeoff times of the takeoff aircraft satisfy that $S_{k_{0}}(\bar{\phi})-S_{k_{0}}(\phi)=S_{k_{1}}(\bar{\phi})-S_{k_{1}}(\phi)=\cdots=S_{k_{s-1}}(\bar{\phi})-S_{k_{s-1}}(\phi)=\mu$ and $S_{k_{s}}(\bar{\phi})=S_{k_{s}}(\phi)$ for some constant $\mu\geq 0$. Suppose that for a given constant $\mu_{0}$, if $0\leq\mu\leq\mu_{0}$, the minimum separation requirements can be satisfied in $\bar{\phi}$ and if $\mu>\mu_{0}$, the minimum separation requirements cannot be satisfied in $\bar{\phi}$. Then the constant $\mu_{0}$ is said to be the initial redundant time of the T-block subsequence $\phi$.
@@
(2) Suppose that $Tcf_{i_{1}-1}\in\phi_{1}$, $Tcf_{i_{1}}\in\phi_{0}$, for the sequence $\phi_{1}$, there is a path from its last aircraft $Tcf_{i_{1}-1}$ to its first aircraft, and for the sequence $\phi_{0}$, there is a path from the second last aircraft to the first aircraft in $\phi_{0}$. If $Tcf_{i_{1}}$ is relevant to only $Tcf_{i_{1}-1}$, it is said that the sequence $\phi$ is a D-block subsequence of $\Phi$, and $(|\phi_{1}|-1)/(|\phi_{0}|-1)$, $S_{i_{1}}(\phi)-S_{i_{0}}(\phi)$ and $S_{i_{1}}(\phi)-S_{i_{2}}(\phi)-Y_{i_{2}i_{1}}$ are the capacity, the time length and the landing time increment of D-block subsequence, where $Tcf_{i_{2}}$ is the second last landing aircraft in $\phi$. Let $\langle Tcf_{k_{0}},Tcf_{k_{1}},\cdots,Tcf_{k_{s}}\rangle$ denote the landing sequence in the D-block subsequence $\phi$. Consider a subsequence $\bar{\phi}$ such that the takeoff times of the takeoff aircraft in $\bar{\phi}$ are consistent with those in $\phi$ and $S_{k_{0}}(\bar{\phi})-S_{k_{0}}(\phi)=S_{k_{1}}(\bar{\phi})-S_{k_{1}}(\phi)=\cdots=S_{k_{s-1}}(\bar{\phi})-S_{k_{s-1}}(\phi)=\mu$ and $S_{k_{s}}(\bar{\phi})=S_{k_{s}}(\phi)$ for some constant $\mu\geq 0$. Suppose that for a given constant $\mu_{0}$, if $0\leq\mu\leq\mu_{0}$, the minimum separation requirements can be satisfied in $\bar{\phi}$ and if $\mu>\mu_{0}$, the minimum separation requirements cannot be satisfied in $\bar{\phi}$. Then the constant $\mu_{0}$ is said to be the initial redundant time of the D-block subsequence $\phi$.
@@
From Tables I-IV, it can be seen that once the separation times between the landing/takeoff aircraft are fixed, in most of the scenarios, each of the takeoff/landing aircraft is relevant to an aircraft with different operation task, which form block subsequences of capacity $1/(*)$. There are only a very small number of scenarios marked in red such that each of the takeoff/landing aircraft is relevant to only an aircraft with the same operation task and block subsequences are not formed. For example, consider an optimal matching problem between two sequences, $\phi_{0}=\langle Tcf_{1},Tcf_{2},Tcf_{3}\rangle$ and $\phi_{1}=\langle Tcf_{4},Tcf_{5},Tcf_{6}\rangle$, where $S_{1}(\phi_{0})=S_{4}(\phi_{1})$, $Y_{12}=Y_{23}=90$, $Y_{45}=60$, and $Y_{56}=80$. If we let $S_{12}(\phi_{0})=S_{23}(\phi_{0})=90$, $S_{45}(\phi_{1})$ and $S_{56}(\phi_{1})$ should be taken as $S_{45}(\phi_{1})=S_{56}(\phi_{1})=90$, whereas if we let $S_{45}(\phi_{1})=60$ and $S_{56}(\phi_{1})=80$, $S_{12}(\phi_{0})$ should be taken as $S_{12}(\phi_{0})=140$, and $S_{23}(\phi_{0})$ can be taken as $S_{23}(\phi_{0})=90$ if there are no other aircraft. In this example, it can be seen that the matching between a landing sequence and a takeoff sequence is not very complicated and the number of all possible block subsequences is $2$.
From Tables I-VIII, it can also be seen that the ranges of the time length and the time increments of block subsequences are both not large as well, in particular when the T-block subsequences are considered. As a matter of fact, based on the calculations for more general scenarios including those listed in Tables I-VIII, it is found that if the number of breakpoint aircraft in a block subsequence is fixed, the number of all possible block subsequences of different class combinations and the ranges of the time lengths and the takeoff/landing time increments of block subsequences are all not large under the minimum separation time standards at Heathrow Airport and for the RECAT-EU system in the absence of resident-point aircraft. This property can be fully explored to find the optimal sequence when a landing sequence and a takeoff sequence are matched.